03 Your parents will retire in 18 years. They currently have $250,000, and they think they will need $1 millin at retirement. What annual interest rate must they earn to reach their goal, assuming they dont save any additional funds? 05 You have $42,180.53 in a brokerage account, and you plan to deposit an additional $5,000 at the end of every future year until your account totals $250,000. You expect to earn 12% annually on the account. How many years will it take to reach your goal? 08 You want to buy a car, and a local bank will lend you $20,000. The loan would be fully amortized over 5 years (60 months), and the nominal interest rate would be 12%, with interest paid monthly, what is the monthly loan payment? what is the loans EFF%? 018 Find the future values of the following ordinary annuities. 1. 2. 3. FV of S400 each 6 months for 5 years at a nominal rate of 12%, compounded semiannually FV of $200 each 3 months for 5 years at a nominal rate of 12%, compounded quarterly The annuities described in parts a and b have the same total amount of money paid into them during the 5-year period, and both earn interest at the same nominal rate, yet the annuity in part b earns $101.75 more than the one in part a over the 5 years. Why doés this occur?
Answer #1

Q3. PV = 250,000
FV = 1,000,0000
Number of Year(n) = 18
Rate = (1,000,000/250,000)1/18 -1 = 8.005%

Q5. PV = 42,180.53
Annuity = 5000
Rate = 12%
FV = 42,180.53*(1+12%)n + 5000*((1+12%)n -1)/12% = 250,000
42,180.53*1.12n + 5000*((1.12n -1)/0.12 = 250,000
42180.53*1.12n + 41666.67*1.12n -41666.67 = 250,000
83847.197*1.12n = 291,666.67
1.12n = 3.47855
applying log on both sides we get
log 1.12n = log 3.47855
n = 11 years